# Axion Astrophysics

### v1

@@ 3.2.1 Axion-photon coupling
If one neglects degeneracy effects and the plasma frequency (a good assumption in plasma conditions when the Primakoff process is the dominating axion production mechanism), one can provide a semi-analytical expression for the energy-loss rate per unit mass in axions [2]
@@
where $T$ and $\rho$ are in K and in g cm-3 respectively. The coefficient $Z(\xi^{2})$ is a function of $\xi^{2}\equiv(\kappa_{s}/2T)^{2}$, with $\kappa_{s}$ the Debye-Huckel screening wavenumber. It can be explicitly expressed as an integral over the photon distribution (see Eq. (4.79) in Ref. [45]). Ref. [2] proposed the analytical parametrization
@@ 4 The Sun and other nearby stars
In more recent years, the Sun has become an important laboratory for exotic physics. This progress was facilitated by a deeper and deeper understanding of the Sun in the past decades, particularly thanks to advancements in helioseismology and neutrino observations. The importance of the Sun in the study of light and weakly interacting particles, such as the axion, relies in large part (though, not only) in its potential to thermally produce these particles. Furthermore, being very close to the Earth, the Sun provides one of the most efficient WISP sources, widely used in several terrestrial experiments. The case of axions, which is the focus of our review, will be discussed extensively below. Other important cases include dark photons [205, 206, 207, 208, 209], and chameleons [210, 211, 212, 213]. It is important to remark that the solar potential as laboratory for exotic processes is not limited to light particles (intended as particle with mass not much larger than the solar core temperature). As we shall see, nuclear reactions or nuclear de-excitation processes occurring in the Sun can also efficiently produce particle of mass considerably larger than its temperature. Finally, the Sun has also been used also to test dark matter Weakly Interacting Massive Particles (WIMPs). In this case, particles in the dark matter halo would be trapped in the Sun core and the further annihilation can lead to an observable flux of high-energy photons or neutrinos [214].
Other MS stars would produce axions in much the same way as the Sun does and are, in this sense, other stellar laboratories for such particles. Obviously, due to their much larger distances, they are not quite as useful, especially as axion factories, as the axion flux decreases quadratically with the distance. However, as we shall see, the flux from all MS stars is not that small and may possibly lead to some interesting phenomenology. Moreover, other stars with properties different from the Sun could provide independent and, in some cases, more efficient axion factories than the Sun itself. The case of supergiant stars, whose core temperature may be considerably larger that that of the Sun, could provide a possibly large source of axions, which could be observed if they convert into photons in the Galactic magnetic field field.
@@ 4.1 Standard Solar Model
    The model is constrained to reproduce today’s solar radius, mass, and luminosity. An essential premise of the SSM posits that the proto-Sun experienced an extensively convective phase, resulting in a uniformly composed Sun until the onset of MS burning. The initial composition by mass is conventionally categorized into hydrogen ($X_{\rm ini}$), helium ($Y_{\rm ini}$), and everything else (the metals, denoted $Z_{\rm ini}$), with $X_{\rm ini}+Y_{\rm ini}+Z_{\rm ini}=1$. The relative proportions of metals are derived from a combination of meteoritic and solar photospheric data. The absolute abundance $Z_{\rm ini}$ can be taken from the modern Sun’s surface abundance $Z_{\rm S}$, after corrections for the effects of diffusion over 4.6 Gyr of solar evolution. Finally, $Y_{\rm ini}/X_{\rm ini}$ is adjusted along with $\alpha_{\rm MLT}$, a parameter describing solar mixing, until the model reproduces the modern Sun’s luminosity and radius. The resulting 4He/H mass fraction ratio is typically $0.27\pm 0.01$, contrasting with the Big-Bang value of $0.23\pm 0.01$, showing that the Sun was formed from previously processed material. The metallicity $Z_{\odot}$ is a fundamental diagnostic of the evolutionary history of the Sun. Up to 1998, the state-of-the-art was given by the spectroscopic measurements of Anders and Grevesse (AG89) [216] and Grevesse and Sauval (GS98) [217]. These presented metallicities $Z_{\odot}=0.0202$ and $Z_{\odot}=0.0170$, respectively. Moreover, heavy element mixtures provided by AG89 and GS98 also yielded good agreement with inferences from helioseismology. However, after 1998 new metallicity determinations, lower than the previous ones, produced a tension between models of the seismic sound-speed profile and depth of the convective zone. In particular, the sets of abundances known as AGS05 [218] and AGSS09 [187] report a metallicity of $Z_{\odot}=0.0122$ and $Z_{\odot}=0.0133$, respectively. This tension between models and helioseismolgy became known as the “solar metallicity problem”. Finally, in 2022 a new photospheric abundance determination [219] reestablished the agreement with helioseismology.
@@ 4.5.3 Search of axions in X-ray telescopes
In Ref. [292, 293] it was proposed to search for solar axions by looking for their conversions into X-rays in the Earth’s magnetosphere with a low-Earth-orbit X-ray detector. It was estimated the sensitivity down to $g_{a\gamma}\lesssim 1.0\times 10^{-11}$ GeV-1 for $m_{a}\lesssim 10^{-4}$ eV. Ref. [294] pointed out that a small fraction axions with keV masses, produced by Primakoff process, with speeds below the escape velocity, would accumulate over time, creating a basin of slow-moving axions trapped on bound orbits. These axions could decay to two photons, yielding an observable signature. Limits on the solar basin of axions were set using recent quiescent solar observations made by the NuSTAR X-ray telescope. However, in Ref. [295] it was remarked that strong cosmological constraints can be placed on dark matter axions decaying into photons, severely limiting the parameter space allowed for the solar basin phenomenology.
@@ 5 Globular Clusters
GCs are the building blocks of any kind of galaxy. They are found in spirals (such as the Milky Way or M31), ellipticals (M87) as well as in Dwarfs Spheroidals or irregular galaxies (e.g., the Magellanic Clouds). Hundreds of GCs populate the Milky Way halo and bulge. They are old ($\sim 11-13$ Gyr) and contain a large number of stars bound by the reciprocal gravitational interaction. Typically a GC contains in between $10^{5}$ and $10^{7}$ stars within a radius of a few parsecs. Most of their stars are nearly coeval, even if there exists a growing amount of observational evidences showing that they host multiple stellar populations characterized by different chemical compositions.
@@ 5.1 The tip of the red giant branch
The RGB of a GC is populated by stars with quite similar masses. The precise value of the mass depends on the cluster age and composition. Assuming an age in the range $11-13$ Gyr, the mass of GC RGB stars is $\sim 0.8-0.9$ M. The internal structure of these stars is characterized by distinct layers and conditions, especially as they approach the brightest luminosity (hereinafter RGB tip). As illustrated in Sec. 2.6.4, the core consists mainly of helium, which has been produced through nuclear fusion of hydrogen in earlier stages of the star’s life. Surrounding the helium core is a shell where hydrogen burns into helium trough the CNO cycle. The He left behind by the H burning makes the core mass progressively bigger. Hence, as the core mass increases, both core density and temperature rise up. Noteworthy, the shell H-burning provides the energy that sustains the losses by radiation at the stellar surface. Hence, as the core mass increases, the shell becomes hotter, thinner and more efficient, so that the stellar luminosity increases. In practice, it exists a well-known relation between the core mass and the stellar luminosity for RGB stars [311]. In addition, owing to the intense energy flux generated by the H burning shell, the outer envelope expands and a deep convective instability sets on. If the core contraction is a source of gravitational energy, energy sink processes are also in action near the center. The combination of energy sources (gravitational contraction and nuclear fusion) and sinks (plasma neutrinos), coupled to an efficient heat conduction (by degenerate electrons), shapes the internal temperature profile (see Fig. 4). Then, when the core mass is $\sim 0.5$ M and the maximum temperature attains $\sim 100$ MK, the fusion of He into C, as triggered by the triple-$\alpha$ reaction, starts. Owing to the degenerate equation of state, a thermonuclear runaway takes place (He-flash, see Fig. 4).
@@
If axions are produced in the core of RGB stars, an additional energy loss would contribute to the shaping of the core temperature profile. At a density of $10^{5}$$10^{6}$ g/cm3, as in the core of a bright RGB star, the plasma frequency becomes of the order of the axion energy, a condition for which the Primakoff process is largely suppressed (see, e.g., Refs. [150, 312]). Similarly, due to the Pauli blocking, the Compton scattering on degenerate electrons is unlikely. Instead axions eventually produced by Bremsstrahlung may provide a non-negligible contribution to the energy-loss [161, 313, 312]. Hence, in case of axions, a higher core mass would be required to attain the conditions for the He ignition and, according to the core mass-luminosity relation, the RGB tip should appear brighter than expected in case of standard energy-loss. In this framework, the RGB-tip luminosity may be used to constrain the strength of the axion-electron coupling $g_{ae}$. In Fig. 27, the energy loss due to Bremsstrahlung for axions is compared to the one of plasma neutrinos, in the core of a star close to the RGB tip, while in Fig. 28, we show the expected variation of the bolometric magnitude [$M_{\rm bol}=4.75-2.5\log(L/L_{\odot})$] of the RGB tip versus the axion-electron coupling strength. Note that the axion production rate scales as $g_{ae}^{2}$ and $T^{4}$. At variance with neutrinos, whose production is heavily concentrated near the center, the axion production peaks at the off-center temperature maximum.
@@
Figure 29: 10 years of $g_{ae}$ bounds from observations of RGB-tip luminosities of galactic GCs: V13 [313], S18 [314], CR20 [315], S20 [312]. The last and most stringent bound (S20+updated distances) has been obtained by adopting the distances reported in [316]. The red arrows represent the corrections required to account for the electron screening (V13) and the $\omega$ Cen distance (CR20) issues (see text for details).
@@
Later on, Ref. [315] brought this bound down to $g_{ae}<1.6\times 10^{-13}$. They used the same RGB stellar models as in Ref. [313], but applying a shift to the predicted RGB tip by-hand. $M_{I}$ to account for the electron-screening issue. This bound, however, was obtained for the cluster $\omega$ Cen (or NGC 5139), whose distance where derived from Ref. [321] by fitting the velocity dispersion profile, after the DR2 Gaia (Global Astrometric Interferometer for Astrophysics) data release, with N-body models of globular clusters. As later noted by Ref. [316], this kinematic distance is about 200 pc shorter than that derived from the parallax and other methods, such as MS fitting or RR Lyrae period-luminosity relation. This discrepancy is likely due to the flattening of $\omega$ Cen ($e>0.17$ [322]). In practices, in the kinematic method, the distance is varied until the observed velocity dispersion profile best matches that of the GC model, but the adopted N-body models are spherically symmetric, so that, in case of a not fully relaxed cluster dynamics, the resulting kinematic distances underestimates the true cluster distance. Summarizing, Ref. [315] adopted the kinematic distance to $\omega$ Cen previously reported by Ref. [321], i.e., $5.24\pm 0.05$ kpc, while Ref. [316] corrected this distance in $5.426\pm 0.047$. However, Ref. [323], based on the parallax from Gaia early data release 3, got a distance of $5.24\pm 0.11$, which is similar to that adopted by Ref. [315], but with a doubled error. In any case, the $g_{ae}$ bound should be relaxed.
@@
A different approach was followed by Ref. [312], where global analysis has been performed of 22 GCs by combining optical and near-infrared data from public archives of ground and space based telescopes. This approach allows one to increase the statistical stellar sample, thus reducing both systematic and statistical errors. In addition, the authors derived cluster distances by scaling the ZAHB luminosity to that of 47 Tuc (NGC 104), for which a rather precise parallax was available after Ref. [324]. Firstly, they applied the RGB-tip method to three GCs, namely, 47 Tuc, M5 and NGC 362, finding $g_{ae}/10^{13}$ bounds of 1.87, 2.30 and 1.37, respectively (95 % C.L.). In addition, they performed a global likelihood analysis combining the RGB-tip measurements of 22 GC, obtaining the more stringent bound to the axion-electron coupling, namely: $g_{ae}<1.48\times 10^{-13}$ (see Fig. 30).
@@
As noted by many authors, among the uncertainties affecting the derived bound, a major contribution to the error budget is given by the distance determination. Also the bolometric corrections (see, e.g., Refs. [325, 326]) used to transform the measured magnitudes into the absolute stellar luminosity (or viceversa) limit the derived constraint to the axion couplings. Concerning distances, an important improvement has been achieved after the third Gaia data release (Gaia-DR3). A quantification of the impact of improved distances can be obtained by adopting the updated distances reported by Ref. [316] for 21 of the 22 GCs of the Ref. [312] sample. In this way an improved bound of $g_{ae}<0.95\times 10^{-13}$ is obtained (see Fig. 29 and Fig. 30).
More in general, since the discrepancy between the standard theoretical prediction and the observed tip luminosity is very small, if any, the derived bound is essentially determined by the error budget. For this reason, statistical and systematic errors affecting both the measured luminosity should be carefully estimated, as well as all the possible uncertainties affecting the theoretical prediction. In any case, even if a further reduction of this global uncertainty will be achieved in the future, when $g_{ae}$ is substantially smaller than $10^{-13}$, the RGB tip becomes practically insensitive to the thermal production of axions, so that other astronomical constraints should be considered to investigate the weak coupling region of the axions parameter space.

### v2

@@ 3.2.1 Axion-photon coupling
If one neglects degeneracy effects and the plasma frequency (a good assumption in plasma conditions when the Primakoff process is the dominating axion production mechanism), one can provide a semi-analytical expression for the energy-loss rate per unit mass in axions [2, 151]
@@
where $T$ and $\rho$ are in K and in g cm-3 respectively. [^11] Note that Ref. [2] has a typo concerning the numerical value of the constant in Eq. (67). The correct value of the constant factor, reported above, can be found in Ref. [151]. The coefficient $Z(\xi^{2})$ is a function of $\xi^{2}\equiv(\kappa_{s}/2T)^{2}$, with $\kappa_{s}$ the Debye-Huckel screening wavenumber. It can be explicitly expressed as an integral over the photon distribution (see Eq. (4.79) in Ref. [45]). Ref. [2] proposed the analytical parametrization
@@ 4 The Sun and other nearby stars
In more recent years, the Sun has become an important laboratory for exotic physics. This progress was facilitated by a deeper and deeper understanding of the Sun in the past decades, particularly thanks to advancements in helioseismology and neutrino observations. The importance of the Sun in the study of light and weakly interacting particles, such as the axion, relies in large part (though, not only) on its potential to thermally produce these particles. Furthermore, being very close to the Earth, the Sun provides one of the most efficient WISP sources, widely used in several terrestrial experiments. The case of axions, which is the focus of our review, will be discussed extensively below. Other important cases include dark photons [207, 208, 209, 210, 211], and chameleons [212, 213, 214, 215]. It is important to remark that the solar potential as laboratory for exotic processes is not limited to light particles (understood as particle with mass not much larger than the solar core temperature). As we shall see, nuclear reactions or nuclear de-excitation processes occurring in the Sun can also efficiently produce particle of mass considerably larger than its temperature. Finally, the Sun has also been used to test dark matter Weakly Interacting Massive Particles (WIMPs). In this case, particles in the dark matter halo would be trapped in the Sun core and the further annihilation can lead to an observable flux of high-energy photons or neutrinos [216].
Other MS stars would produce axions in much the same way as the Sun does and are, in this sense, other stellar laboratories for such particles. Obviously, due to their much larger distances, they are not quite as useful, especially as axion factories, as the axion flux decreases quadratically with the distance. However, as we shall see, the flux from all MS stars is not that small and may possibly lead to some interesting phenomenology. Moreover, other stars with properties different from the Sun could provide independent and, in some cases, more efficient axion factories than the Sun itself. The case of supergiant stars, whose core temperature may be considerably larger than that of the Sun, could provide a possibly large source of axions, which could be observed if they convert into photons in the Galactic magnetic field.
@@ 4.1 Standard Solar Model
    The model is constrained to reproduce today’s solar radius, mass, and luminosity. An essential premise of the SSM posits that the proto-Sun experienced an extensively convective phase, resulting in a uniformly composed Sun until the onset of MS burning. The initial composition by mass is conventionally categorized into hydrogen ($X_{\rm ini}$), helium ($Y_{\rm ini}$), and everything else (the metals, denoted $Z_{\rm ini}$), with $X_{\rm ini}+Y_{\rm ini}+Z_{\rm ini}=1$. The relative proportions of metals are derived from a combination of meteoritic and solar photospheric data. The absolute abundance $Z_{\rm ini}$ can be taken from the modern Sun’s surface abundance $Z_{\rm S}$, after corrections for the effects of diffusion over 4.6 Gyr of solar evolution. Finally, $Y_{\rm ini}/X_{\rm ini}$ is adjusted along with $\alpha_{\rm MLT}$, a parameter describing solar mixing, until the model reproduces the modern Sun’s luminosity and radius. The resulting 4He/H mass fraction ratio is typically $0.27\pm 0.01$, contrasting with the Big-Bang value of $0.23\pm 0.01$, showing that the Sun was formed from previously processed material. The metallicity $Z_{\odot}$ is a fundamental diagnostic of the evolutionary history of the Sun. Up to 1998, the state-of-the-art was given by the spectroscopic measurements of Anders and Grevesse (AG89) [218] and Grevesse and Sauval (GS98) [219]. These presented metallicities $Z_{\odot}=0.0202$ and $Z_{\odot}=0.0170$, respectively. Moreover, heavy element mixtures provided by AG89 and GS98 also yielded good agreement with inferences from helioseismology. However, after 1998 new metallicity determinations, lower than the previous ones, produced a tension between models of the seismic sound-speed profile and depth of the convective zone. In particular, the sets of abundances known as AGS05 [220] and AGSS09 [189] report a metallicity of $Z_{\odot}=0.0122$ and $Z_{\odot}=0.0133$, respectively. This tension between models and helioseismolgy became known as the “solar metallicity problem” (see also Ref. [221] for exotic physics implications for this problem). Finally, in 2022 a new photospheric abundance determination [222] reestablished the agreement with helioseismology.
@@ 4.5.3 Search of axions in X-ray telescopes
In Ref. [295, 296] it was proposed to search for solar axions by looking for their conversions into X-rays in the Earth’s magnetosphere with a low-Earth-orbit X-ray detector. It was estimated the sensitivity down to $g_{a\gamma}\lesssim 1.0\times 10^{-11}$ GeV-1 for $m_{a}\lesssim 10^{-4}$ eV. Ref. [297] pointed out that a small fraction axions with keV masses, produced by Primakoff process, with speeds below the escape velocity, would accumulate over time, creating a basin of slow-moving axions trapped on bound orbits. These axions could decay to two photons, yielding an observable signature. Limits on the solar basin of axions were set using recent quiescent solar observations made by the NuSTAR X-ray telescope. However, in Ref. [298] it was remarked that strong constraints can be placed on axions produced in the early universe, decaying into photons, severely limiting the parameter space allowed for the solar basin phenomenology.
@@ 5 Globular Clusters
GCs are the building blocks of any kind of galaxy. They are found in spirals (such as the Milky Way or M31), ellipticals (M87) as well as in Dwarf Spheroidals or irregular galaxies (e.g., the Magellanic Clouds). Hundreds of GCs populate the Milky Way halo and bulge. They are old ($\sim 11-13$ Gyr) and contain a large number of stars bound by the reciprocal gravitational interaction. Typically a GC contains in between $10^{5}$ and $10^{7}$ stars within a radius of a few parsecs. Most of their stars are nearly coeval, even though there exists a growing amount of observational evidences showing that they host multiple stellar populations characterized by different chemical compositions.
@@ 5.1 The tip of the red giant branch
The RGB of a GC is populated by stars with quite similar masses. The precise value of the mass depends on the cluster age and composition. Assuming an age in the range $11-13$ Gyr, the mass of GC RGB stars is $\sim 0.8-0.9$ M. The internal structure of these stars is characterized by distinct layers and conditions, especially as they approach the brightest luminosity (hereinafter RGB tip). As illustrated in Sec. 2.6.4, the core consists mainly of helium, which has been produced through nuclear fusion of hydrogen in earlier stages of the star’s life. Surrounding the helium core is a shell where hydrogen burns into helium through the CNO cycle. The He left behind by the H burning makes the core mass progressively bigger. Hence, as the core mass increases, both core density and temperature rise up. Noteworthy, the shell H-burning provides the energy that sustains the losses by radiation at the stellar surface. Hence, as the core mass increases, the shell becomes hotter, thinner and more efficient, so that the stellar luminosity increases. In practice, it exists a well-known relation between the core mass and the stellar luminosity for RGB stars [314]. In addition, owing to the intense energy flux generated by the H burning shell, the outer envelope expands and a deep convective instability sets on. If the core contraction is a source of gravitational energy, energy sink processes are also in action near the center. The combination of energy sources (gravitational contraction and nuclear fusion) and sinks (plasma neutrinos), coupled to an efficient heat conduction (by degenerate electrons), shapes the internal temperature profile (see Fig. 4). Then, when the core mass is $\sim 0.5$ M and the maximum temperature attains $\sim 100$ MK, the fusion of He into C, as triggered by the triple-$\alpha$ reaction, starts. Owing to the degenerate equation of state, a thermonuclear runaway takes place (He-flash, see Fig. 4).
@@
If axions are produced in the core of RGB stars, an additional energy loss would contribute to the shaping of the core temperature profile. At a density of $10^{5}$$10^{6}$ g/cm3, as in the core of a bright RGB star, the plasma frequency becomes of the order of the axion energy, a condition for which the Primakoff process is largely suppressed (see, e.g., Refs. [152, 315]). Similarly, due to the Pauli blocking, the Compton scattering on degenerate electrons is unlikely. Instead axions eventually produced by Bremsstrahlung may provide a non-negligible contribution to the energy-loss [163, 316, 315]. Hence, If Bremsstrahlung axions are actually produced in RGB stars, a higher core mass would be required to attain the conditions for the He ignition and, according to the core mass-luminosity relation, the RGB tip should appear brighter than expected in case of standard energy-loss. In this framework, the RGB-tip luminosity may be used to constrain the strength of the axion-electron coupling $g_{ae}$. In Fig. 27, the energy loss due to Bremsstrahlung axions is compared to the one of plasma neutrinos, in the core of a star close to the RGB tip, while in Fig. 28, we show the expected variation of the bolometric magnitude [$M_{\rm bol}=4.75-2.5\log(L/L_{\odot})$] of the RGB tip versus the axion-electron coupling strength. Note that the axion production rate scales as $g_{ae}^{2}$ and $T^{4}$. Unlike neutrinos, whose production is heavily concentrated near the center, the axion production peaks at the off-center temperature maximum.
@@
Figure 29: 10 years of $g_{ae}$ bounds from observations of RGB-tip luminosities of galactic GCs: V13 [316], S18 [317], CR20 [318], S20 [315], T24 [319]. The S20(+updated distances) has been obtained by adopting the distances reported in [320]. The two arrows represent the corrections required to account for the electron screening (V13) and the $\omega$ Cen distance (CR20) issues (see text for details).
@@
Later on, Ref. [318] brought this bound down to $g_{ae}<1.3\times 10^{-13}$. The authors used the same RGB stellar models as in Ref. [316], but applying a shift to the predicted $I$ magnitude of the RGB tip to account for the electron-screening issue. This bound was obtained from the RGB tip of the cluster $\omega$ Cen (or NGC 5139), whose distance was derived from Ref. [325] by fitting the velocity dispersion profile, after the second data release of the Global Astrometric Interferometer for Astrophysics (Gaia-DR2), with N-body models of globular clusters. As later noted by Ref. [320], this kinematic distance is about 200 pc shorter than that derived from the parallax and other methods, such as MS fitting or RR Lyrae period-luminosity relation. This discrepancy is likely due to the flattening of $\omega$ Cen ($e>0.17$ [326]). In practice, in the kinematic method, the distance is varied until the observed velocity dispersion profile best matches that of spherically-symmetric N-body models. Therefore, in case of a not fully relaxed cluster dynamics, the resulting kinematic distance underestimates the true cluster distance. Summarizing, Ref. [318] adopted the kinematic distance to $\omega$ Cen previously reported by Ref. [325], i.e., $5.24\pm 0.05$ kpc, while Ref. [320] corrected this distance in $5.426\pm 0.047$. However, Ref. [327], based on the parallax from Gaia early data release 3, got a distance of $5.24\pm 0.11$, which is similar to that adopted by Ref. [318], but with a doubled error. In any case, the $g_{ae}$ bound should be relaxed. Noteworthy, Ref. [318] also derive a new bound $g_{ae}<1.6\times 10^{-13}$ from the RGB tip of NGC 4258, a spiral galaxy in the Canes Venatici constellation. Owing to the presence of an active galactic nucleus (AGN), NGC 4258 is also classified as a Seyfert 2 galaxy with a characteristic H2O maser emission from the accretion disk around the central black hole. Such a maser source allows an accurate distance determination for this galacy Ref. [328]. However, unlike GCs, a galaxy hosts many generations of stars, characterized by different ages and chemical compositions. In practice, the superposition of the giant stars belonging to different stellar generations implies a spread in both the $I$ magnitude and the $V-I$ color of the stars close to the RGB tip. Thus, to account for these variations, a correction is applied to the observed $I$ magnitude of the brightest RGB stars, as derived from an empirical $I$-$(V-I)$ relation (see, e.g., Ref. [329]). In this way, a unique absolute $I$ magnitude ($M^{TRGB}_{I}$) is provided at the same reference color, i.e., $(V-I)^{TRGB}=1.8$ mag. The choice of this reference color is not arbitrary. Indeed, it coincides approximately to the color of the brightest stars in the GC M5, thus allowing the use of the same theoretical models adopted by [316], eventually corrected to account for the electron-screening issue. For this reason, [318] also applied the same correction to the $M^{TRGB}_{I}$ of the GC $\omega$ Cen, a procedure that would not have been necessary if they had used stellar models with composition and age appropriate for this stellar cluster. In any case, the use of this relation introduces a further uncertainty to the derived bound, which is probably small only in the case of a limited spread in the $V-I$ color. In this context, a controversial conclusion was reached by [330]181818 This paper has been uploaded to the ArXiv on May 4, 2023 and, so far, never accepted for publication in a peer-reviewed journal.. Basing on a detailed statistical analysis, they argued that the correlations among the uncertainties of only 3 model parameters, i.e, metallicity, He abundance and age, imply a substantially weaker than previously found bound for the axion-electron coupling. Similarly to [318], they compared the observed $I$ magnitude of the RGB tip of one globular cluster ($\omega$ Cen) and two galaxies (NGC 4258 and the Large Magellanic Cloud) with those predicted by a machine learning (ML) code, previously trained with a large set of RGB stellar models. However, they assumed an unrealistic large range of values for the 3 input parameters, ignoring the many spectroscopic and photometric constraints. Moreover, as in [318], they applied a correction to the observed $I$ magnitude to account for the differences between the observed and the reference $V-I$ colors. As previously noted, this correction is needed in case of a galaxy showing a superposition of multiple stellar populations characterized by different ages and compositions, but it is unnecessary in the case of a globular cluster whose composition and age have been accurately measured. Then, at variance with [318], they also apply the same correction to the theoretical prediction. In practice, the predicted $M^{TRGB}_{I}$ given by the ML is converted to a prediction at the reference $V-I$ color. This correction, when coupled to the errors of the cluster parameters, amplify the uncertainty on the predicted $M^{TRGB}_{I}$. For example, a variation of the metallicity within its error bar, implies a change of the theoretical $V-I$ and, in turn, a different correction to the predicted $I$ magnitude. In this way, the many uncertainties affecting the the theoretical $V-I$ are propagated into the predicted $M^{TRGB}_{I}$. In addition to the adopted bolometric corrections, the resulting axion-electron coupling bound also depends on the model uncertainties due to the radiative opacity, the surface boundary conditions, the equation of state of partially ionized matter, as well as to the adopted mixing-length parameter and the mass-loss rate. Summarizing, [330] demonstrated that the method they adopted is heavily affected by these model uncertainties and, for this reason, it fails to produce a reliable and robust constraint for the axion-electron coupling. Note that the previous bounds obtained by [316, 318] are not affected by this theoretical $V-I$ uncertainties, except for those related to the bolometric corrections.
@@
A different approach was followed by Ref. [315] (also see [317]), where 22 GCs have been considered by combining optical and near-infrared data from public archives of ground and space based telescopes. Thanks to these multi-wavelength analyses, they derived the bolometric magnitude of the brightest RGB stars of each cluster, a quantity that can be directly compared to the corresponding luminosity predicted by stellar models accounting for the additional axion cooling. Distances were derived by scaling the ZAHB luminosity to that of 47 Tuc (NGC 104), for which a rather precise parallax was available after Ref. [331]. As it is well known, the luminosity of a bright RGB star does not depends on its total mass, but, rather, on its He-core mass [314]. Therefore, the uncertainties affecting the physics of the H-rich envelope, e.g., those related to the mixing-length parameter, the mass-loss rate, the atomic and molecular opacity or the adopted surface boundary conditions, produce negligible effects on the predicted luminosity and, in turn, on the predicted bolometric magnitude. Note that this is a major advantage of this method with respect to the procedure followed in Refs. [316, 318, 330]. First, they applied this method to three GCs, namely, 47 Tuc, M5 and NGC 362, finding $g_{ae}/10^{13}$ bounds of 1.87, 2.30 and 1.37, respectively (95 % C.L.). Moreover, they performed a global likelihood analysis combining the RGB-tip measurements of 22 GC, obtaining the more stringent bound to the axion-electron coupling, namely: $g_{ae}<1.48\times 10^{-13}$ (see Fig. 30). This approach allows to reduce both systematic and statistical errors. The authors also analyzed the various uncertainties of the stellar models and how they affect the predicted RGB-tip luminosity. Also, comparisons with results obtained by means of other stellar evolution codes were discussed in detail.
@@
As noted by many authors, among the uncertainties affecting the derived bound, a major contribution to the error budget is given by the distance determination. Also the bolometric corrections (see, e.g., Refs. [332, 333]) used to transform the measured magnitudes into the absolute stellar luminosity (or viceversa) limit the derived constraint to the axion couplings. Concerning distances, an important improvement has been achieved after the third Gaia data release (Gaia-DR3). A quantification of the impact of improved distances can be obtained by adopting the updated distances reported by Ref. [320] for 21 of the 22 GCs of the Ref. [315] sample. In this way an improved bound of $g_{ae}<0.95\times 10^{-13}$ may be obtained (see Fig. 29 and Fig. 30). While we were finishing writing this review, a new analysis based on Gaia-DR3 photometry and astrometry of GCs was uploaded to the ArXiv by Ref. [319] [^19] This paper has been recently accepted for publication on Pis’ma v ZhETF.. The derived bound, i.e., $g_{ae}<0.58\times 10^{-13}$ (95% CL), confirms and reinforces previous results.
More in general, since the discrepancy between the standard theoretical prediction and the observed tip luminosity is very small, if any, the derived bound is essentially determined by the error budget. For this reason, statistical and systematic errors affecting the measured stellar luminosity should be carefully estimated, as well as all the possible uncertainties affecting the corresponding theoretical predictions. In any case, even if a further reduction of this global uncertainty will be achieved in the future, when $g_{ae}$ is substantially smaller than $10^{-13}$, the RGB tip becomes practically insensitive to the thermal production of axions, so that other astronomical constraints should be considered to investigate the extremely weak coupling region of the axion parameter space.